### v11

# Causal Bayesian Optimization via Exogenous Distribution Learning

Abstract
Maximizing a target variable as an operational objective in a structural causal model is an important problem. Causal Bayesian Optimization (CBO) methods either rely on interventions that alter the causal structure to maximize the reward; or introduce action nodes to endogenous variables so that the data generation mechanisms are adjusted to achieve the objective. This paper introduces a novel method to learn the distribution of exogenous variables, which is typically marginalized through expectation or ignored by existing CBO methods. Exogenous distribution learning improves the approximation accuracy of structural causal models in a surrogate model that is usually trained with limited observational data. Moreover, the learned exogenous distribution extends existing CBO to general causal schemes beyond simple Additive Noise Models (ANMs). The recovery of exogenous variables allows us to use a more flexible prior for noise or unobserved hidden variables. We develop a new CBO method by leveraging the learned exogenous distribution. Experiments on different datasets and applications show the benefits of our proposed method.
1 Introduction
Bayesian Optimization (BO) has broad applications, such as automating modern industrial processes, drug discovery, and more general synthetic biology, that require optimizing black-box objective functions [15; 3; 10]. In many real-world problems, we have some structural knowledge of the unknown functions that can be leveraged to improve BO efficiency. CBO has been developed to utilize the structural information of the objective [1; 2; 22]. CBO combines ideas from causal inference, uncertainty quantification, and sequential decision-making. It generalizes Bayesian Optimization, which treats the input variables of the objective function as independent, to scenarios where causal information between input variables is available [1]. CBO has been developed for optimizing medical and ecological interventions as illustrative examples [1; 2].
1.1 Approach and Contributions
In this paper, we develop an EXogenous distribution learning augmented Causal Bayesian Optimization method, EXCBO. Given observation data points of a structural causal model (SCM, [18; 17]), the exogenous variable associated with each endogenous node is recovered with an encoder-decoder framework. The learned distribution of the exogenous variable is then approximated with the density of the recovered exogenous variable that is parameterized with a flexible model, e.g. Gaussian Mixture. The recovered exogenous distribution improves the surrogate model’s accuracy in the SCM approximation. Different from existing CBO methods [1; 2; 22] that typically focus on Additive Noise Models (ANMs, [11]), our method enables us to extend CBO to general causal models beyond ANMs. Our flexible surrogate model has the benefit of better causal inference regarding CBO updating.
The contributions in this paper include
  • We propose a method to recover the exogenous noise variable of each endogenous node in an SCM using observational data; the exogenous distribution in each node is then learned using a flexible model that is capable of modeling multi-mode density functions;
  • The flexibility of learning exogenous distributions naturally extends the proposed CBO framework to general causal models beyond ANMs.
  • We provide a theoretical study on the recovery of exogenous variables. Moreover, we give a theoretical analysis of the algorithm regret bound. Our analysis shows that the proposed EXCBO algorithm achieves sublinear cumulative regret bounds.
  • Experiments have been performed to validate the benefits of exogenous distribution learning and the proposed EXCBO algorithm. Applications on epidemic dynamic model calibration and COVID-19 testing are discussed to support the proposed CBO framework.
In the following parts of this paper, Section 2 presents the background and related work; Section 3 sets up the problem studied in this paper with the introduction of the proposed CBO framework; Section 4 presents the exogenous recovery method; The proposed SCM algorithm, EXCBO, is detailed in Section 5 with theoretical analysis of EXCBO given in Section 6; Section 7 presents experimental results; Section 8 lists additional remarks; Section 9 concludes the paper.
2 Background
In this section, we give a brief review of SCM, intervention, and CBO methods.
2.1 Structural Causal Model
An SCM is denoted by $\mathcal{M}=(\mathcal{G},\mathbf{F},\mathbf{V},\mathbf{U})$. Here $\mathcal{G}$ is a directed acyclic graph (DAG), and $\mathbf{F}=\{f_{i}\}_{i=0}^{d}$ stands for $d+1$ mechanisms of all variables. $\mathbf{V}$ is the endogenous variable set, and $\mathbf{U}$ is the exogenous (background) variable set. Samples of the $i$th endogenous variable are generated with
$X_{i}=f_{i}(\mathbf{Z}_{i},U_{i});\ \mathbf{Z}_{i}=\mathbf{pa}(i),\ U_{i}\sim p(U_{i}),\ \text{for}\ i\in[d].$ (1)
Here $d]=\{0,1,...,d\}$. We use $X_{i}$ to represent the $i$th variable as well as the $i$th node in $\mathcal{G}$, and $\mathbf{pa}(i)$ stands for the parents of $i$, and $\mathbf{ch}(i)$ is the children of node $X_{i}$ regarding $\mathcal{G}$. In addition, we have $U_{i}\perp\!\!\!\perp\mathbf{Z}_{i}$ and $U_{i}\perp\!\!\!\perp U_{j},\forall i\neq j$. $f_{i}$ is a mapping from $\mathbb{R}^{|\mathbf{pa}(i)|+1}\rightarrow\mathbb{R}$. $\mathcal{X}_{i}$, $\mathbf{\mathcal{Z}}_{i}$, and $\mathcal{U}_{i}$ stand for the value domain of $X_{i}$, $\mathbf{Z}_{i}$, and $U_{i}$, respectively. We further assume that the expectation of $X_{i},i\in[d]$, i.e. $\mathbb{E}[X_{i}]$, exists. Existing CBO methods [1; 22] usually assume that the exogenous variable is an ANM [11] to the data generation mechanism, i.e., $X_{i}=f_{i}(\mathbf{Z}_{i})+U_{i},\ \text{with}\ U_{i}\sim\mathcal{N}(0,1).$
2.2 Intervention
In SCM $\mathcal{M}$, let $\mathbf{I}$ be a subset of endogenous variables, i.e. $\mathbf{I}\subset\mathbf{U}$. The intervention on target $\mathbf{I}$ is given by $\mathbf{F}_{x}=\{f_{i}|X_{i}\notin\mathbf{I}\}\cup\{f_{j}|X_{j}\in\mathbf{I}\}.$ The intervention on variable $X_{j}$ is implemented with the change of the data generation mechanism from $f_{i}$ to a constant scalar. There are generally two types of interventions: hard and soft intervention. The hard intervention on target $\mathbf{I}$ is given by $\mathbf{F}_{x}=\{f_{i}|X_{i}\notin\mathbf{I}\}\cup\{f_{j}:=\alpha_{j}|X_{j}\in\mathbf{I}\}$. Here $\mathbf{\alpha}$ is the realization of variables in $\mathbf{I}$. The do-operation [18] $do(\mathbf{X}_{\mathbf{I}}:=\bm{\alpha})$ represents the hard intervention that changes the SCM from $\mathcal{M}$ to $\mathcal{M}_{\mathbf{\alpha}}$. A hard intervention on a node breaks its causal connection with its parents.
This paper focuses on the soft intervention (or imperfect intervention, [19]). Following Model-based CBO (MCBO, [22]), we introduce an action variable to each endogenous variable, i.e., $\mathbf{F}_{x}=\{f_{i}|X_{i}\notin\mathbf{I}\}\cup\{f_{j}:=f_{j}(\mathbf{Z}_{j},A_{j},U_{j})|X_{j}\in\mathbf{I}\}.$ Here $\mathbf{Z}_{j}=\mathbf{pa}(j)$. With the soft intervention on target $\mathbf{I}$, the data generation mechanism becomes
$X_{i}=\begin{cases}f_{i}(\mathbf{Z}_{i},U_{i}),&\quad\text{if }X_{i}\notin\mathbf{I}\\ f_{i}(\mathbf{Z}_{i},A_{i},U_{i}),&\quad\text{if }X_{i}\in\mathbf{I}\end{cases},$ (2)
where $A_{i}$ is the continuous action variable for $X_{i}$ if $X_{i}\in\mathbf{I}$. Action $A_{i}$ with the corresponding value domain $\mathcal{A}_{i}$ controls the change of the data generation mechanism of node $X_{i}$. The soft intervention can be represented with Pearl’s do-operation [18] $do\big(\mathbf{X}_{\mathbf{I}}:=\mathbf{f}(\mathbf{Z}_{\mathbf{I}},A,U_{\mathbf{I}})\big)$ .
2.3 Function Network Bayesian Optimization
Similar to CBO, the structure of the functions in Function Network BO (FNBO, [4; 3]) is assumed to be given, and the parameters and specific form of each function are unknown. FNBO operates soft interventions and uses an expected improvement (EI) acquisition function to select actions. FNBO assumes that the system is noiseless, which could be a restrictive assumption in practice. Both FNBO and CBO belong to a broad research direction using observations of a function network to improve the sample efficiency of vanilla BO algorithms [5].
2.4 Causal Bayesian Optimization
In CBO, a sequence of actions is performed to interact with an SCM $\mathcal{M}$. The graph structure, DAG $\mathcal{G}$ in $\mathcal{M}$ is given, with the edge functions $\mathbf{F}=\{f_{i}\}_{i=0}^{d}$ being fixed but unknown. Probabilistic surrogate models (typically Gaussian Processes, i.e. GPs, [23]) are implemented to search for the intervention actions in attaining the optimal objective/reward value. In [1], the authors proposed a CBO algorithm that obtains both the optimal intervention target set and the corresponding input values to reach the optimal objective value in an SCM. A dynamic CBO (DCBO) algorithm is proposed in [2], where all causal effects in the SCM $\mathcal{M}$ are changing over time.
In MCBO [22], soft interventions are optimized to achieve the maximum target variable of an SCM. With action variables, the edge function set becomes $\mathbf{F}=\{f_{i}\}_{i=0}^{d}$ with $f_{i}:\mathcal{Z}_{i}\times\mathcal{A}_{i}\to\mathcal{X}_{i}$. Let $x_{i,t}$ denote the observation of node $X_{i}$ at time step $t$, and $i\in[d]$, $t\in[T]$. $T$ is the total number of time steps. At step $t$ MCBO selects actions $a_{:t}=\{a_{i,t}\}_{i=0}^{d}$ and obtains observations $\mathbf{x}_{:,t}=\{x_{i,t}\}_{i=0}^{d}$. The relation between the action $a_{i,t}$ and the observation is given by an additive noise model, i.e., $x_{i,t}=f_{i}(\mathbf{z}_{i,t},a_{i,t})+u_{i,t},\ \ \forall i\in[d].$ At the target node $d$, $a_{d,t}=0$, and $y_{t}=f_{d}(\mathbf{z}_{d,t},a_{d,t})+u_{d,t},$ and here $y_{t}$ depends on the whole intervention vector. The action vector that maximizes the expected reward regarding the exogenous variable is $\mathbf{a}^{*}$, i.e. $\mathbf{a}^{*}=\operatorname*{arg\,max}_{\mathbf{a}\in\mathcal{A}}\mathbb{E}[y|\mathbf{a}].$ The GP surrogate model is implemented to learn the reward function and guide BO in optimizing $y$.
3 Problem Statement
Following existing CBO algorithms [1; 2; 22], the DAG $\mathcal{G}$ is assumed to be known in this paper. Similar to MCBO [22], our framework adopts GP surrogate models to guide the optimization of soft interventions controlled by an action vector $\mathbf{a}=\{a_{i}\}_{i=0}^{d}$ in reaching the optimal reward value. This section gives details about the problem settings studied in this paper.
3.1 Motivations for Exogenous Distribution Learning
In existing CBO models, exogenous distributions are ignored or marginalized to achieve the simplicity of interventions [1; 2; 22]. As mentioned in previous sections, we propose an encoder-decoder scheme to recover the exogenous variable of each endogenous node in an SCM. The distribution of an exogenous variable ($U_{i}$) is approximated with the density of the recovered surrogate exogenous variable ($\widehat{U}_{i}$) which is modeled with a flexible model, e.g. Gaussian Mixture. The recovered exogenous distribution improves the surrogate model in the approximation of the SCM. Thus different from existing methods [1; 2; 22] that typically rely on the ANM assumption, our EXCBO allows us to extend CBO to a broader scope of causal models beyond ANMs. In addition, with a more accurate surrogate model, our method could achieve improved reward and objective values.
3.2 Decomposable Generation Mechanism
In our problem, we assume the DAG structure of the SCM $\mathcal{M}$, i.e. $\mathcal{G}$, is given, and the edges or the mechanisms of the endogenous variables, i.e. $\mathbf{F}=\{f_{i}\}_{i=0}^{d}$, are fixed but unknown. We further assume that the problem is causal sufficient, which means that there are no unobserved endogenous variables or confounders. We present the definition of decomposable generation mechanism (DGM) used in this paper as follows.
Definition 3.1.
(DGM) The data generation function $f$ follows decomposable generation mechanism (DGM) if $X=f(\mathbf{Z},U)=f_{a}(\mathbf{Z})+f_{b}(\mathbf{Z})f_{c}(U)$, and the mappings are $f_{a}():\mathcal{Z}\to\mathbb{R}$, $f_{b}():\mathcal{Z}\to\mathbb{R}$, and $f_{c}():\mathcal{U}\to\mathbb{R}$.
In a DGM, $f_{c}()$ can be any continuous function, including nonlinear and non-monotonic one-dimensional function classes regarding $U$. $f_{b}(\mathbf{Z})$ can be any continuous function with $f_{b}(\mathbf{Z})\neq 0$. The last term of a DGM, i.e. $f_{b}(\mathbf{Z})f_{c}(U)$, indicates that given a fixed value of the parent $Z$, the generated variable $X$’s variance is not only determined by $U$ but also by the value of $\mathbf{Z}$. Therefore, with $f_{b}(\mathbf{Z})$ DGM represents a large class of mechanisms that both $U$ and $\mathbf{Z}$ affect the variance of the generated variable $X$. DGM is different from the LSNMs (location-scale noise models, [12]), which adopts a locally linear relationship between the exogenous variable $U$ and the endogenous variable $X$.
We show that the distributions of exogenous variables can be recovered if the data generation mechanism $f$ in each node (1) belongs to DGM in Section 4.1.
3.3 CBO via Exogenous Distribution Learning
Unlike ANMs in existing CBO algorithms, e.g. MCBO [22] (briefly reviewed in Section 2.4), we here propose to flexibly model the mapping $f_{i}()$ with exogenous variables, for which we propose a new EXCBO framework for CBO through exogenous distribution learning.
Figure 1: EXCBO: Causual Bayesian Optimization through exogenous distribution learning. The distribution of $U_{i}$ is approximated with the density of the recovered surrogate $\widehat{U}_{i}$. The EXCBO algorithm searches for the action vector $a$ that maximizes the reward $Y$.
At step $t$ of CBO, the relation between actions $\mathbf{a}_{i,t}$, $\mathbf{u}_{i,t}$ and the observation $\mathbf{x}_{i,t}$ is given by
$\mathbf{x}_{i,t}=\mathbf{f}_{i}(\mathbf{z}_{i,t},\mathbf{a}_{i,t},\mathbf{u}_{i,t})\ \ \forall i\in[d].$ (3)
Our surrogate model in EXCBO takes the exogenous variable $U_{i}$ as the input as shown in (3) and Figure 1. Let $\mathcal{R}$ represent the set of the root nodes. As the parent of root nodes is null, we set $\mathbf{z}_{i,t}=\mathbf{0}$, if $i\in\mathcal{R}$. Similarly, we have $a_{d,t}=0$ at the target node $d$,
$y_{t}=f_{d}(\mathbf{z}_{d,t},a_{d,t},u_{d,t}).$
Given an action vector $\mathbf{a}=\{a_{i}\}_{i=0}^{d}$ and $\mathbf{u}=\{u_{i}\}_{i=0}^{d}$, the reward or target value is simply denoted with $y=\mathbf{F}(\mathbf{a},\mathbf{u})$. The problem becomes searching for an action vector maximizing the objective:
$\mathbf{a}^{*}=\operatorname*{arg\,max}_{\mathbf{a}\in\mathcal{A}}\mathbb{E}[y|\mathbf{a}],$ (4)
where the expectation is taken over $u$. The goal of our method is to design a sequence of interventions $\{\mathbf{a}_{t}\}_{t=0}^{T}$ that achieves a high average expected reward. In Section 4, we will show that if a mechanism $f_{i}$} in $\mathbf{F}=\{f_{i}\}_{i=0}^{d}$ meets the DGM assumption, the exogenous variable $U_{i}$ and its distribution can be learned from observations. With the learned distributions of all the exogenous variables, each action’s reward can be evaluated in the operations of equation (4).
To analyze the convergence of our method, we study the cumulative regret over a time horizon of $T$:
$R_{T}=\sum_{t=1}^{T}\bigg[\mathbb{E}[y|\mathbf{a}^{*}]-\mathbb{E}[y|a_{:,t}]\bigg].$
Cumulative regret bound helps us to analyze the convergence property of the algorithm. In the experiments, we use the attained objective value or reward value $y$ as the metric to compare our method and existing algorithms. A better metric depends on the real-world application and the effectiveness of the optimized action sequence.
4 Exogenous Distribution Learning
Given the observations of an endogenous node and its parents in a SCM, our method is to recover the distribution of the node’s exogenous variable. The learning of the exogenous distribution is implemented with GPs. We first focus on the exogenous distribution learning regarding one node.
4.1 Exogenous Variable Recovery for One Node
According to (2), an endogenous variable $X_{i}$ either has an action variable $A_{i}$ or not. To better illustrate the idea with compact notations, we use $\mathbf{Z}_{i}$ to represent variable $X_{i}$’s parents and its action variable in this section (Section 4), i.e., $\mathbf{Z}_{i}=(\mathbf{Z}_{i},A_{i})$ if $X_{i}\in\mathbf{I}$. Hence the learning of node $X$’s exogenous distortion becomes a problem of recovering $U$’s distribution given the observations of $X$ and $\mathbf{Z}$ regarding equation $X=f(\mathbf{Z},U)$. To make our presentation concise, we have the following definition for the causal mechanism of the triplets $(\mathbf{Z},U,X)$ that appears as a node in an SCM.
Definition 4.1.
($\tau$-SCM) Let $X$ and $U$ be one-dimensional variables, and $f()$ is the causal mechanism generating $X$ from variables $Z$ and $U$, i.e. $X=f(\mathbf{Z},U)$. We assume $\mathbf{Z}\perp\!\!\!\perp U$, and $(\mathbf{Z},U,X,f)$ is defined as a $\tau$-SCM.
[…]
In a $\tau$-SCM, $\mathbf{Z}$ is the parent of $X$, and it may be multi-dimensional; $U$ is the exogenous variable of $X$, and it is different from the bijective generation mechanisms (BGM, [16]). In BGM with data generation mechanism $X=f(\mathbf{Z},U)$, $f()$ is required to be monotonic and invertible regarding $X$ and $U$ given a value of $\mathbf{Z}$. The encoder-decoder framework (shown in Figure 2) is used to recover a surrogate of the exogenous variable.
Definition 4.2.
[…]
Theorem 4.1.
Let $(\mathbf{Z},U,X,f)$ be a $\tau$-SCM, and $(\widehat{U},\phi,h,g)$ is an EDS surrogate of the exogenous variable $U$. We assume that $f()$ is differentiable and has a decomposable structure as $X=f(\mathbf{Z},U)=f_{a}(\mathbf{Z})+f_{b}(\mathbf{Z})f_{c}(U)$, and $f_{b}(\mathbf{Z})\neq 0$. Then with a constant $a$, $\widehat{U}=a\big(f_{c}(U)-\mathbb{E}[f_{c}(U)]\big)$, $\mathbb{E}[\widehat{U}]=0$, and $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$.
Figure 3: Scopes of different mechanisms.
We use the distribution of $\widehat{U}=\mathbf{s}(U)=h(\mathbf{Z},X)$, i.e. p(U^)p($\widehat${U}), to represent $p(U)$ in the surrogate model. In our surrogate model, $f()$ is approximated with the learned $g()$ and $\widehat{u}$, i.e., $x=f(\mathbf{z},u)=g(\mathbf{z},\widehat{u})=g(\mathbf{z},\mathbf{s}(u))$. Figure 3 gives the relationship between different mechanisms. These features substantially extend the identification of exogenous variable $U$ to much broader nonlinear and non-monotonic data-generation mechanism classes compared to ANM (linear) and BGM (monotonic, [16]). DGM is significantly different from the ANM [11] mechanism given by $X=f(\mathbf{pa}(X))+U$, which is a linear monotonic function of $U$ and widely used by many existing BO and CBO methods. We use EDS to represent the $\tau$-SCMs with the exogenous distribution that can be approximately recovered by using EDS (with or without the condition $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$) and Figure 3 gives the scopes of different mechanisms including EDS.
4.2 Implementation of Exogenous Distribution Learning
There are multiple ways to implement the encoder-decoder framework in Figure 2, e.g., via Variational Auto-Encoder (VAE) [13]. To make the implementation simple, we implement both the encoder and decoder using GP regression models based on EDS definition 4.2. For a node with an action variable $A$, the decoder becomes $g():\mathcal{Z}\times\mathcal{A}\times\widehat{\mathcal{U}}\to\mathcal{X}$, the encoder becomes $h():\mathcal{Z}\times\mathcal{A}\times\mathcal{X}\to\widehat{\mathcal{U}}$, and $\phi():\mathcal{Z}\times\mathcal{A}\to\mathcal{X}$. Here $g()$ and $\phi()$ are implemented with GP regression models [23].
We use a Gaussian Mixture model to fit the recovered distribution of $\widehat{U}$, i.e. p(U^)p($\widehat${U}). We use p(U^)p($\widehat${U}) to replace $p(U)$ in the computing of the probabilistic surrogate objective. For the nodes in SCM $\mathcal{M}$, let $\mathbf{G}=\{g_{i}\}_{i=0}^{d}$, and $\mathbf{H}=\{h_{i}\}_{i=0}^{d}$.
5 CBO with Exogenous Distribution Learning
In this section, we present the EXCBO algorithm, describing the probabilistic model and acquisition function used.
5.1 Statistical Model
In our model, the function $f_{i}$ that generates variable $X_{i}$ is learned through $g_{i}$, and $X_{i}=g_{i}(\mathbf{Z}_{i},A_{i},\widehat{U}_{i})$. We use GPs [23] to learn the surrogate of $g_{i}$, i.e., $\tilde{g}_{i}$. For $i\in[d]$, let $\mu_{g,i,0}$ and $\sigma_{g,i,0}$ denote the prior mean and variance function for each $f_{i}$, respectively.
Following [8], at time $t$, let $\tilde{\mathbf{G}}$ be the statistically plausible function set of $\mathbf{G}$, i.e., $\tilde{\mathbf{G}}=\{\tilde{g}_{i}\}_{i=0}^{d}$. Similarly, the plausible model of $\mathbf{H}$ is denoted by $\tilde{\mathbf{H}}=\{\tilde{h}_{i}\}_{i=0}^{d}$. Moreover, at step $t$, the observation set is $\mathcal{D}_{t}=\{\mathbf{z}_{:,1:t},a_{:,1:t},x_{:,1:t}\}$. The posterior of $g_{i}$ with the input of node $i$, $(\tilde{\mathbf{z}}_{i},\tilde{a}_{i},\tilde{\widehat{u}}_{i})$, is given by
$\tilde{g}_{i,t}(\tilde{\mathbf{z}}_{i},\tilde{a}_{i},\tilde{\widehat{u}}_{i})\sim\mathcal{GP}(\mu_{g,i,t-1},\sigma^{2}_{g,i,t-1});$
$\mu_{g,i,t-1}=\mu_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{a}_{i},\tilde{\widehat{u}}_{i});\quad\sigma_{g,i,t-1}=\sigma_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{a}_{i},\tilde{\widehat{u}}_{i}).$
Then $\tilde{x}_{i,t}=\tilde{g}_{i,t}(\tilde{\mathbf{z}}_{i},\tilde{a}_{i},\tilde{\widehat{u}}_{i})$ denotes observations from one of the plausible models. Here $\tilde{\widehat{u}}_{i}\sim p(\widehat{U}_{i})$ in the sampling of the learned distribution of $\widehat{U}_{i}$.
Given an observation $(\mathbf{z}_{i},a_{i},x_{i})$ at node $i$, the exogenous recovery $\widehat{u}_{i}=h_{i}(\mathbf{z}_{i},a_{i},x_{i})=\frac{x_{i}-\mu_{\phi,i}(\mathbf{z}_{i},a_{i})}{\sigma_{\phi,i}(\mathbf{z}_{i},a_{i})}$. At time step $t$, the posterior of $\phi_{i}$ with the input of node $i$, $(\mathbf{z}_{i},a_{i})$, is given by
$\tilde{\phi}_{i,t}(\mathbf{z}_{i},a_{i})\sim\mathcal{GP}\big(\mu_{\phi,i,t-1}(\mathbf{z}_{i},a_{i}),\sigma^{2}_{\phi,i,t-1}(\mathbf{z}_{i},a_{i})\big)$ (5)
Therefore, $\tilde{\widehat{u}}_{i}=\tilde{h}_{i,t}(\mathbf{z}_{i},a_{i},x_{i})=\frac{x_{i}-\mu_{\phi,i,t-1}(\mathbf{z}_{i},a_{i})}{\sigma_{\phi,i,t-1}(\mathbf{z}_{i},a_{i})}$. According to the definition of $h()$ in Theorem 4.1, $h()$ also follows a GP, i.e. $h_{i,t}(\mathbf{z}_{i},a_{i},x_{i})\sim\mathcal{GP}(\mu_{h,i,t-1},\sigma^{2}_{h,i,t-1}).$ This GP is defined by $\tilde{\phi}_{i,t}()$ which is sampled with (5). Different from $g_{i}()$, the observations of the input $(\mathbf{Z}_{i},A_{i},X_{i})$ for $h_{i}()$ are only required at the training time, and we only need to sample the learned $p(\widehat{U}_{i})$ to get value $\widehat{u}_{i}$ for model prediction or model sampling.
5.2 Acquisition Function
Algorithm 1 describes the proposed method solving (4). In iteration $t$, it uses GP posterior belief of $y$ to construct an upper confidence bound (UCB, [7]) of $y$:
$\text{UCB}_{t-1}(\mathbf{a})=\mu_{t-1}(\mathbf{a})+\beta_{t}\sigma_{t-1}(\mathbf{a}).$ (6)
Here
$\mu_{t-1}(\mathbf{a})=\mathbb{E}[\mu_{g,d,t-1}(\tilde{\mathbf{z}}_{d},\tilde{a}_{d},\tilde{\widehat{u}}_{d})]\ ;\ \ \sigma_{t-1}(\mathbf{a})=\mathbb{E}[\sigma_{g,d,t-1}(\tilde{\mathbf{z}}_{d},\tilde{a}_{d},\tilde{\widehat{u}}_{d})],$
where the expectation is taken over p(U^)p($\widehat${U}). In (6), $\beta_{t}$ controls the tradeoff between exploration and exploitation of Algorithm 1. The UCB-based algorithm is a classic strategy that is widely used in BO and stochastic bandits [14; 20]. The proposed algorithm adapts the “optimism in the face of uncertainty” (OFU) strategy by taking the expectation of the UCB as part of the acquisition process.
5.3 Algorithm
Let $k_{g,i},k_{\phi,i}$, $\forall i\in[d]$ represent the kernel functions of $g_{i}$ and $\phi_{i}$. The proposed EXCBO algorithm is summarized by Algorithm 1. In each iteration, a new sample is observed according to the UCB values. Then the posteriors of $\mathbf{G}$ and $\mathbf{H}$ are updated with the new dataset. The next section gives a theoretical analysis of the algorithm.
Algorithm 1 EXCBO
Input: $k_{g,i},k_{\phi,i}$, $\forall i\in[d]$
Result: Intervention actions $a_{i}$, $\forall i\in[d]$
for $t=1$ to $T$ do
  Find $\mathbf{a}_{t}$ by optimizing the acquisition function, $\mathbf{a}_{t}\in\operatorname*{arg\,max}\text{UCB}_{t-1}(\mathbf{a})$;
  Observe samples $\{\mathbf{z}_{i,t},x_{i,t}\}_{i=0}^{d}$ with the action sequence $\mathbf{a}_{t}$ and update $\mathcal{D}_{t}$;
  Use $\mathcal{D}_{t}$ to update posteriors $\{\mu_{\phi,i,t},\sigma^{2}_{\phi,i,t}\}_{i=0}^{d}$ and sample the exogenous surrogate $\{\widehat{u}_{i,t}\}_{i=0}^{d}$;
  Use $\mathcal{D}_{t}\cup\{\widehat{u}_{i,t}\}_{i=0}^{d}$ to update the decoder posteriors $\{\mu_{g,i,t},\sigma^{2}_{g,i,t}\}_{i=0}^{d}$ ;
end for
6 Theoretical Study
This section describes the convergence guarantees for EXCBO using soft interventions. Our analysis follows the theoretical study in [22] and shows that EXCBO has a sublinear cumulative regret bound.
In DAG $\mathcal{G}$ over $\{X_{i}\}_{i=0}^{d}$, let $N$ be the maximum distance from a root to $X_{d}$, i.e., $N=\max_{i}\text{dist}(X_{i},X_{d})$. Here $\text{dist}(\cdot,\cdot)$ is a measure of the edges in the longest path from $X_{i}$ to the reword node $X_{d}$. Let $M$ denote the maximum number of parents of any variables in $\mathcal{G},M=\max_{i}|\mathbf{pa}(i)|$. Let $L_{t}$ be a function of $L_{g}$, $L_{\sigma_{g}}$, and $N$. With Assumptions B.1B.3 in the Appendix, the following theorem bounds the performance of EXCBO in terms of cumulative regret.
We present the assumptions used in the regret analysis in the Appendix. Assumption B.1 gives the Lipschitz conditions of $g_{i}$, $\sigma_{g,i}$, and $\mu_{g,i}$. It holds if the RKHS of each $g_{i}$ has a Lipschitz continuous kernel [9; 22]. Assumption B.3 holds when we assume that the $i$th GP prior uses the same kernel as the RKHS of $g_{i}$ and that $\beta_{i,t}$ is sufficiently large to ensure the confidence bounds in
$\bigg|g_{i}(\mathbf{z}_{i},a_{i},\widehat{u}_{i})-\mu_{g,i,t-1}(\mathbf{z}_{i},a_{i},\widehat{u}_{i})\bigg|\leq\beta_{i,t}\sigma_{g,i,t-1}(\mathbf{z}_{i},a_{i},\widehat{u}_{i})\ ,\ \ \ \forall\mathbf{z}_{i}\in\mathcal{Z}_{i},a_{i}\in\mathcal{A}_{i},\widehat{u}_{i}\in\widehat{\mathcal{U}}_{i}.$
Theorem 6.1.
Consider the optimization problem in (4), with the SCM satisfying Assumptions B.1B.3, where $\mathcal{G}$ is known but $\mathbf{F}$ is unknown. Then with probability at least $1-\alpha$, the cumulative regret of Algorithm 1 is bounded by
$R_{T}$ $\leq\mathcal{O}(L_{T}M^{N}d\sqrt{T\gamma_{T}}).$
Here $\gamma_{T}=\max_{t}\gamma_{i,T}$ denote the maximum information gain at time $T$. This maximum information gain is commonly used in many Bayesian Optimizations [20]. Many common kernels, such as linear and squared exponential kernels, lead to sublinear information gain in $T$, and it results in an overall sublinear regret for EXCBO [22]. The proof of Theorem 6.1 and further analysis can be found in the Appendix.
The analysis in this paper focuses on the DGM mechanisms. To extend the analysis to BGMs, we need to consider the computation cost involving the independence penalization on variables $\widehat{U}$ and $\mathbf{Z}$. For mechanisms beyond DGMs and BGMs, we conjecture that the surrogate approximation accuracy may decrease, but the convergence rate may not decrease a lot.
7 Experimental Study
This section presents experimental comparisons of the proposed EXCBO and existing algorithms.
7.1 Baselines
We compare EXCBO against three algorithms: UCB, EICF, and MCBO, where UCB is a conventional method of BO [7]; EICF [3] is the composite function method of BO; MCBO [22] is the CBO method discussed in previous sections. Unlike other methods, MCBO relies on neural networks along with GPs to model uncertainty. All algorithms are implemented with Python and BoTorch [6], and we use the implementations of UCB, EICF, and MCBO in the package published with [22; 21]. The initial observation size for MCBO is the default value from the package [21], which is $2(|A|+1)$, and $|A|$ is the number of action variables. For other methods, the initial sample size is in the range of $[5,20]$. We use four different random seeds to run all the algorithms to obtain the means and standard derivations of the reward values. All the experiments are conducted on an HPC cluster with CentOS 7 OS, 48-core nodes, and 10 GB of requested memory.
7.2 Alpine2 Dataset
Figure 4: Alpine2 graph structure.
We study the algorithms using the Alpine2 dataset. There are six endogenous nodes in the Alpine2 dataset as shown in Figure 4. The exogenous distributions for $X$ and $Y$ are Gaussian Mixture models with two components, i.e.,
$p(U)=w_{1}\mathcal{N}(\mu_{1},c_{1}\sigma)+w_{2}\mathcal{N}(\mu_{2},c_{2}\sigma),\ \ \,\,w_{1},w_{2},c_{1},c_{2}>0.$ (7)
(a) $\sigma=0.05$                    (b) $\sigma=0.1$                   (c) $\sigma=0.2$
Refer to caption Refer to caption Refer to caption
Figure 5: Results of Alpine2 at different noise levels.
Due to the excessively long computing time used by MCBO, we have two baselines in this set of experiments, i.e., UCB and EICF. The results of Alpine2 are shown in Figures 5-(a-c). As shown in the plots, our EXCBO performs the best on all three noise levels. It proves the effectiveness and benefits of the proposed EXCBO method.
7.3 Dropwave Dataset
Figure 6: Graph structure of Dropwave dataset.
The structure of Dropwave is given in Figure 6. Figure 7 gives the average reward values achieved by the competing algorithms using four different random seeds in the code, and shadows are the scaled stand derivations of the four results. In each round, the size of the observation on the true model equals the round number.
(a) $\sigma=0.05$                                   (b) $\sigma=0.3$
Refer to caption Refer to caption
Figure 7: Results of Dropwave at different noise levels.
Figure 7-(a) presents the results at $\sigma=0.05$. It can be observed clearly that EXCBO achieves the best reward at nearly all rounds. Figure 7-(b) gives the results at $\sigma=0.3$. In Figure 7-(b), EXCBO exceeds UCB and EICF at round 130, and it achieves the best performance at round 193. It indicates that our EXCBO method may require more training samples to recover the exogenous distribution especially when the variance of the noise is large.
Empirically, the proposed EXCBO uses a similar CPU time as UCB for the same round number on different datasets. EICF takes a little bit more computational cost compared to EXCBO and UCB. However, MCBO consumes much more CPU time ($10+$ times) compared to the other three methods that do not rely on neural networks. This means that EXCBO is more scalable in terms of computational cost in comparison with the existing state-of-the-art methods.
7.4 Epidemic Model Calibration
Figure 8: Graph structure for epidemic model calibration.
We test EXCBO on an epidemic model calibration by following the setup in [4]. In this model, as shown in Figure 8, $I_{i,t}$ represents the fraction of the population in group $i$ that are “infectious” at time $t$; $\beta_{i,j,t}$ is the rate of the people from group $i$ who are “susceptible” have close physical contact with people in group $j$ who are “infectious” at time $t$. We assume there are two groups, and infections resolve at a rate of $\gamma$ per period. The number of infectious individuals in group $i$ at the start of the next time period is $I_{i,t+1}=I_{i,t}(1-\gamma)+(1-I_{i,t})\sum_{j}\beta_{i,j,t}I_{j,t}$. We assume each $I_{i,t}$ has an observation noise $U_{i,t}$. The model calibration problem is that given limited noisy observations of $I_{i,t}$s, how to efficiently find the $\beta_{i,j,t}$ values in the model. The reward is defined as the negative mean square error (MSE) of all the $I_{i,t}$ observations as the objective function to optimize. In this model, $\beta_{i,j,t}$s are the action variables.
(a) $\sigma=0.1$                                  (b) $\sigma=0.3$
Refer to caption Refer to caption
Figure 9: (a-b): Results of epidemic model calibration.
Figure 9-(a-b) visualize the results at the noise levels with $\sigma=0.1$ and $\sigma=0.3$. The noise is added with two models as in (7). As shown in Figure 9-(a-b), our EXCBO always performs better than state-of-the-art model calibration methods in both cases, and our method has a faster convergence rate compared to the baselines. With the capability to recover and learn the exogenous distributions, our method is more robust and stable in this application scenario.
7.5 Pooled Testing for COVID-19
Figure 10: Graph structure for COVID-19 pooled testing problem.
(a) $\sigma=0.2$                                      (b) $\sigma=0.4$
Refer to caption Refer to caption
Figure 11: (a-b): Results of COVID-19 pooled testing problem.
We further compare EXCBO and existing methods using the COVID-19 pooled testing problem [4]. The graphical structure is given by Figure 10. In Figure 10, $I_{t}$ is the fraction of the population that is infectious at time $t$; $R_{t}$ is the fraction of the population that is recovered and cannot be infected again, and time point $t\in\{1,2,3\}$. The additional fraction $S_{t}=1-I_{t}-R_{t}$ of the population is susceptible and can be infected. During each period $t$, the entire population is tested using a pool size of $x_{t}$. The loss $L_{t}$, incorporates the costs resulting from infections, testing resources used, and individuals isolated at period $t$. The objective is to choose pool size $x_{t}$ to minimize the total loss $\sum_{t}L_{t}$. Therefore, $x_{t}$s are the action variables/nodes that the algorithms try to optimize to achieve lower costs. Figure 11-(a-b) gives the optimization results from different methods, where the reward $y=-\sum_{t}L_{t}$. The proposed method EXCBO achieves the best rewards at both noise levels $\sigma=0.2$ and $\sigma=0.4$. MCBO’s inferior results on this dataset might be because of the over-fitting issue of the neural networks. The experimental results on different datasets show that the learned exogenous distributions improve EXCBO’s capability to reach the optimal reward values. The experiments show that the learning of exogenous variable distribution provides a more accurate surrogate model given an SCM and observational data.
8 Additional Remarks
8.1 Limitations
There are two main limitations of this work. Firstly, the proposed exogenous variable recovery method (EDS) only works under the DGM or BGM assumptions. DGMs and BGMs are a small portion of the whole data generation mechanisms. We leave the empirical and theoretical study of EDS on non-DGM and non-BGM mechanisms to future works. Secondly, empirically, our EXCBO method may require more training samples to achieve superior performance when the variance of the noise or exogenous variable increases. It means that datasets with stronger noise or additional uncertainty may require more training samples to recover the exogenous distributions. But eventually, our method could outperform existing methods given enough data samples.
8.2 Impact Statements
As a new causal Bayesian optimization framework, EXCBO may help reduce the required training samples for more efficient and cost-effective decision-making, which may have broader impacts in many science and engineering applications, such as future pandemic preparedness with better-calibrated epidemic dynamic models as illustrated in the paper. However, if misused, the societal consequences of designing new systems or materials with unforeseen future threats has to be taken into consideration with caution.
9 Conclusions
We propose a novel Causal Bayesian Optimization algorithm, EXCBO, that approximately recovers the endogenous variables in a structured causal model. With the recovered exogenous distribution, our method naturally improves the surrogate model’s accuracy in the approximation of the SCM. Furthermore, the recovered exogenous variables may enhance the surrogate model’s capability in causal inference and hence improve the reward values attained by EXCBO. We additionally provide theoretical analysis on both exogenous variable recovery and the algorithm’s cumulative regret bound. Experiments on multiple datasets show the algorithm’s soundness and benefits.
Appendix A Analysis on Exogenous Variable Recovery
A.1 Proof of Theorem 4.1
Before we prove Theorem 4.1, we present a similar result for ANMs [11].
Theorem A.1.
[…]
Proof.
As $\phi(\mathbf{z})$ is an optimal approximation of $\rho(X,\mathbf{z})$, with $\mathbf{Z}\perp\!\!\!\perp U$, we have
$\phi(\mathbf{z})$ $=\mathbb{E}[\rho(X,\mathbf{z})]=\mathbb{E}[\rho_{1}(\mathbf{z})+\rho_{2}(U)]=\int\big(\rho_{1}(\mathbf{z})+\rho_{2}(u)\big)p(u)du$
$=\rho_{1}(\mathbf{z})+\mathbb{E}[\rho_{2}(U)].$
Thus, the decoder becomes
$h(\mathbf{z},x)$ $=\rho(x,\mathbf{z})-\phi(\mathbf{Z}=\mathbf{z})$
$=\rho_{1}(\mathbf{z})+\rho_{2}(u)-\rho_{1}(\mathbf{z})-\mathbb{E}[\rho_{2}(U)]$
$=\rho_{2}(u)-\mathbb{E}[\rho_{2}(U)].$
Therefore, $\widehat{U}=h(\mathbf{Z},X)=\rho_{2}(U)-\mathbb{E}[\rho_{2}(U)]$ is a function of $U$, and $h(\mathbf{Z},X)\perp\!\!\!\perp\mathbf{Z}$, i.e., $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$. ∎
Example A.1.
For an ANM [11] model $X=f(\mathbf{Z})+U$, we have $\rho(X,\mathbf{Z})=X$, $\rho_{1}(\mathbf{Z})=f(\mathbf{Z})$, and $\rho_{2}(U)=U$, then $\widehat{U}=h(\mathbf{Z},X)=U-\bar{U}$ .
Example A.2.
For a model $X=2Ze^{-U}-e^{-Z}$, we have $\rho(X,Z)=\log(X+e^{-Z})$, $\rho_{1}(Z)=\log(2Z)$, and $\rho_{2}(U)=-U$, then $\widehat{U}=h(Z,X)=-U+\bar{U}$ .
Example A.1 shows that the exogenous variable in any ANM model is identifiable. In practice, variable $X$’s generation mechanism $f()$ is generally unknown, and it is hard to propose a general form function $\rho()$ that can perform on any $f()$s and transform them to ANMs.
Theorem 4.1 Let $(\mathbf{Z},U,X,f)$ be a $\tau$-SCM, and $(\widehat{U},\phi,h,g)$ is an EDS surrogate of the exogenous variable $U$. We assume that $f()$ is differentiable and has a decomposable structure as $X=f(\mathbf{Z},U)=f_{a}(\mathbf{Z})+f_{b}(\mathbf{Z})f_{c}(U)$, and $f_{b}(\mathbf{Z})\neq 0$. Then with a constant $a$, $\widehat{U}=a\big(f_{c}(U)-\mathbb{E}[f_{c}(U)]\big)$, $\mathbb{E}[\widehat{U}]=0$, and $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$.
Proof.
$\forall\mathbf{z}\in\mathcal{Z}$, as $\phi(\mathbf{z})$ is an optimal approximation of any value of $X=f(\mathbf{z},U)$, with $\mathbf{Z}\perp\!\!\!\perp U$, we have the mean function as
$\mu_{\phi}(\mathbf{z})$ $=\mathbb{E}[X(\mathbf{z},U)]=\int\big(f_{a}(\mathbf{z})+f_{b}(\mathbf{z})f_{c}(u)\big)p(u)du$
$=f_{a}(\mathbf{z})+\int f_{b}(\mathbf{z})f_{c}(u)p(u)du$
$=f_{a}(\mathbf{z})+f_{b}(\mathbf{z})\mathbb{E}\big[f_{c}(U)\big].$
Then with $U=u$,
$x-\mu_{\phi}(\mathbf{z})$
$=$ $f_{a}(\mathbf{z})+f_{b}(\mathbf{z})f_{c}(u)-f_{a}(\mathbf{z})-f_{b}(\mathbf{z})\mathbb{E}\big[f_{c}(U)\big]$
$=$ $f_{b}(\mathbf{z})\big(f_{c}(u)-\mathbb{E}[f_{c}(U)]\big).$ (8)
With (8), the variance of regression model $\phi()$ is
$\mathbb{E}[(X-\mu_{\phi}(\mathbf{z}))^{2}]$
$=$ $\mathbb{E}\bigg[f_{b}(\mathbf{z})\big(f_{c}(U)-\mathbb{E}[f_{c}(U)]\big)^{2}f_{b}(\mathbf{z})\bigg]$
$=$ $f^{2}_{b}(\mathbf{z})\mathbb{E}\bigg[\big(f_{c}(U)-\mathbb{E}[f_{c}(U)]\big)^{2}\bigg]$
$=$ $f^{2}_{b}(\mathbf{z})\sigma_{f_{c}}^{2}.$
As $\mathbf{Z}\perp\!\!\!\perp U$, as a function of $\mathbf{Z}$, the learned variance $\sigma^{2}_{\phi}(\mathbf{z})$ does not capture the information of $U$. $\sigma^{2}_{\phi}(\mathbf{z})$ learns the variance function with respect to variable $\mathbf{Z}$, i.e., $f^{2}_{2}(\mathbf{z})$. Therefore, $\sigma_{\phi}(\mathbf{z})=c|f_{b}(\mathbf{z})|$. Then, by (8),
$\frac{x-\mu_{\phi}(\mathbf{z})}{\sigma_{\phi}(\mathbf{z})}=$ $\frac{f_{b}(\mathbf{z})\big(f_{c}(u)-\mathbb{E}[f_{c}(U)]\big)}{c|f_{b}(\mathbf{z})|}$
$=$ $\frac{s}{c}\big(f_{c}(u)-\mathbb{E}[f_{c}(U)]\big).$ (9)
Here $s=\textrm{sign}[f_{b}(\mathbf{z})]$. Let $a=\frac{s}{c}$, we define
$\widehat{U}:=\frac{X-\mu_{\phi}(\mathbf{Z})}{\sigma_{\phi}(\mathbf{Z})}=a\big(f_{c}(U)-\mathbb{E}[f_{c}(U)]\big).$
It shows that $\mathbb{E}[\widehat{U}]=0$, and $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$. ∎
Appendix B Regret Analysis
Our analysis follows the study in [22]. In the DAG $\mathcal{G}$ over $\{X_{i}\}_{0}^{d}$, let $N$ be the maximum distance from a root to $X_{d}$, i.e., $N=\max_{i}\text{dist}(X_{i},X_{d})$. Here $\text{dist}(\cdot,\cdot)$ is a measure of the edges in the longest path from $X_{i}$ to the reword node $X_{d}$. Let $M$ denote the maximum number of parents of any variables in $\mathcal{G},M=\max_{i}|\mathbf{pa}(i)|$. Let $L_{t}$ be a function of $L_{g}$, $L_{\sigma_{g}}$. According to Theorem 4.1, with the EDS structure given in Figure 2 in the main text, the exogenous variable and its distribution can be recovered. For each observation of the dynamic surrogate model, we assume the sampling of p(U^)p($\widehat${U}), $\tilde{\widehat{u}}=\mathbf{s}(\tilde{u})=\mathbf{s}(u)$.
We give the assumptions used in the regret analysis. Assumption B.1 gives the Lipschitz conditions of $g_{i}$, $\sigma_{g,i}$, and $\mu_{g,i}$. It holds if the RKHS of each $g_{i}$ has a Lipschitz continuous kernel [9; 22]. Assumption B.3 holds when we assume that the $i$th GP prior uses the same kernel as the RKHS of $g_{i}$ and that $\beta_{i,t}$ is sufficiently large to ensure the confidence bounds in
$\bigg|g_{i}(\mathbf{z}_{i},a_{i},\widehat{u}_{i})-\mu_{g,i,t-1}(\mathbf{z}_{i},a_{i},\widehat{u}_{i})\bigg|\leq\beta_{i,t}\sigma_{g,i,t-1}(\mathbf{z}_{i},a_{i},\widehat{u}_{i})\ ,\ \ \ \forall\mathbf{z}_{i}\in\mathcal{Z}_{i},a_{i}\in\mathcal{A}_{i},\widehat{u}_{i}\in\widehat{\mathcal{U}}_{i}.$
Assumption B.1.
$\forall g_{i}\in\mathbf{G}$, $g_{i}$ is $L_{g}$-Lipschitz continuous; moreover, $\forall i,t$, $\mu_{g,i,t}$ and $\sigma_{g,i,t}$ are $L_{\mu_{g}}$ and $L_{\sigma_{g}}$ Lipschitz continuous.
Assumption B.2.
$\forall f_{i}\in\mathbf{F}$, $f_{i}$ is differentiable and has a decomposable structure with $X=f_{i}(\mathbf{Z}_{i},U_{i})=f_{i(a)}(\mathbf{Z}_{i})+f_{i(b)}(\mathbf{Z}_{i})f_{i(c)}(U_{i})$, and $f_{i(b)}(\mathbf{z}_{i})\neq 0,\forall\mathbf{z}_{i}\in\mathcal{Z}_{i}$.
Assumption B.3.
$\forall i,t$, there exists sequence $\beta_{i,t}\in\mathbb{R}_{>0}$, with probability at least $(1-\alpha)$, for all $\mathbf{z}_{i},a_{i},\widehat{u}_{i}\in\mathcal{Z}_{i}\times\mathcal{A}_{i}\times\widehat{\mathcal{U}}_{i}$ we have $\big|g_{i}(\mathbf{z}_{i},a_{i},\widehat{u}_{i})-\mu_{g,i,t-1}(\mathbf{z}_{i},a_{i},\widehat{u}_{i})\big|\leq\beta_{i,t}\sigma_{g,i,t-1}(\mathbf{z}_{i},a_{i},\widehat{u}_{i})$, and $|h(\mathbf{z}_{i},a_{i},x_{i})-\mu_{h,i,t-1}(\mathbf{z}_{i},a_{i},x_{i})|\leq\beta_{i,t}\sigma_{h,i,t-1}(\mathbf{z}_{i},a_{i},x_{i})$.
The following lemma bounds the value of $\tilde{\widehat{u}}$ with the variance of the encoder.
Lemma B.1.
$\|\widehat{u}_{i,t}-\tilde{\widehat{u}}_{i,t}\|\leq 2\beta_{t}\|\sigma_{\widehat{u}_{i,t-1}}\|=2\beta_{t}\|\sigma_{h,i,t-1}\|.$
Proof.
With Assumption B.3 and $\widehat{u}_{i,t}=h_{i,i-1}(\mathbf{z}_{i},a_{i},x_{i})$, let $\tilde{\widehat{u}}_{i,t}=\mu_{\widehat{u}_{i,t-1}}\mathbf{z}_{i},a_{i},x_{i}+\beta_{t}\sigma_{\widehat{u}_{i,t-1}}(\mathbf{z}_{i},a_{i},x_{i})\circ\bm{\omega}_{\widehat{u}_{i,t-1}}(\mathbf{z}_{i},a_{i},x_{i})$, and here $|\bm{\omega}_{\widehat{u}_{i,t-1}}(\mathbf{z}_{i},a_{i},x_{i})|\leq 1$. Then
$\|\widehat{u}_{i,t}-\tilde{\widehat{u}}_{i,t}\|=$ $\|\tilde{\widehat{u}}_{i,t}-\mu_{\widehat{u}_{i,t-1}}(\mathbf{z}_{i},a_{i},x_{i})-\beta_{t}\sigma_{\widehat{u}_{i,t-1}}(\mathbf{z}_{i},a_{i},x_{i})\circ\bm{\omega}_{\widehat{u}_{i,t-1}}(\mathbf{z}_{i},a_{i},x_{i})\|$
$\leq$ $\|\tilde{\widehat{u}}_{i,t}-\mu_{\widehat{u}_{i,t-1}}(\mathbf{z}_{i},a_{i},x_{i})\|+\beta_{t}\|\sigma_{\widehat{u}_{i,t-1}}(\mathbf{z}_{i},a_{i},x_{i})\circ\bm{\omega}_{\widehat{u}_{i,t-1}}(\mathbf{z}_{i},a_{i},x_{i})\|$
$\leq$ $2\beta_{t}\|\sigma_{\widehat{u}_{i,t-1}}(\mathbf{z}_{i},a_{i},x_{i})\|=2\beta_{t}\|\sigma_{h,i,t-1}\|.$
With the decomposable Assumption B.2 on $f_{i}$, $\sigma^{2}_{h,i,t-1}\propto f^{2}_{i(b)}(\mathbf{z}_{i},a_{i})\big(f_{i(c)}(U)-\mathbb{E}[f_{i(c)}(U)]\big)^{2}$ according to the proof of Theorem 4.1. $f_{i(b)}()$ is learned with the variance of regression model $\phi()$, i.e. $\sigma_{\phi,i,t}()$.
Lemma B.2.
$\|x_{d,t}-\tilde{x}_{d,t}\|\leq 2\beta_{t}M^{N_{i}}(2\beta_{t}L_{\sigma_{g}}+L_{g})^{N_{i}}\sum_{j=0}^{i}\big(\sigma_{g,j,t-1}(\mathbf{z}_{j,t})+\sigma_{\widehat{u}_{j,t-1}}\big).$
Proof.
We use $g_{i}(\mathbf{z}_{i,t},\widehat{u}_{i,t})$ to represent $g_{i}(\mathbf{z}_{i,t},a_{i,t},\widehat{u}_{i,t})$ because we assume the actions to be the same for the process generating $x_{i,t}$ and $\tilde{x}_{i,t}$. Similarly, $\mu_{g,i,t-1}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})=\mu_{g,i,t-1}(\tilde{\mathbf{z}}_{i,t},\tilde{a}_{i,t},\tilde{\widehat{u}}_{i,t})$, $\sigma_{g,i,t-1}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})=\sigma_{g,i,t-1}(\tilde{\mathbf{z}}_{i,t},\tilde{a}_{i,t},\tilde{\widehat{u}}_{i,t})$.
We use the reparameterization trick, and write $\tilde{x}_{i,t}$ as
$\tilde{x}_{i,t}=\tilde{g}_{i}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})=\mu_{g,i,t-1}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})+\beta_{t}\sigma_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})\circ\bm{\omega}_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t}).$
Here $|\bm{\omega}_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})|\leq 1$. Hence, we have
$\|x_{i,t}-\tilde{x}_{i,t}\|$ $=\|g_{i}(\mathbf{z}_{i,t},\widehat{u}_{i,t})-\mu_{g,i,t-1}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})-\beta_{t}\sigma_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})\bm{\omega}_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})\|$
$=\|g_{i}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})-\mu_{g,i,t-1}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})-\beta_{t}\sigma_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})\bm{\omega}_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})$
$~~~~~~~+g_{i}(\mathbf{z}_{i,t},\widehat{u}_{i,t})-g_{i}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})\|$
$\leq\|g_{i}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})-\mu_{g,i,t-1}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})\|+\|\beta_{t}\sigma_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})\bm{\omega}_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})\|$
$~~~~~~~~~+\|g_{i}(\mathbf{z}_{i,t},\widehat{u}_{i,t})-g_{i}(\tilde{\mathbf{z}}_{i,t},\tilde{\widehat{u}}_{i,t})\|$
$\overset{\zeta_{1}}{\leq}\beta_{t}\|\sigma_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})\|+\beta_{t}\|\sigma_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})\|+L_{g_{i}}\big\|[\mathbf{z}_{i,t};\widehat{u}_{i,t}]-[\tilde{\mathbf{z}}_{i,t};\tilde{\widehat{u}}_{i,t}]\big\|$
$=2\beta_{t}\|\sigma_{g,i,t-1}(\mathbf{z}_{i},\widehat{u}_{i,t})+\sigma_{g,i,t-1}(\tilde{\mathbf{z}}_{i},\tilde{\widehat{u}}_{i,t})-\sigma_{g,i,t-1}(\mathbf{z}_{i},\widehat{u}_{i,t})\|+L_{g_{i}}\big\|[\mathbf{z}_{i,t};\widehat{u}_{i,t}]-[\tilde{\mathbf{z}}_{i,t};\tilde{\widehat{u}}_{i,t}]\big\|$
$\overset{\zeta_{2}}{\leq}2\beta_{t}\bigg(\|\sigma_{g,i,t-1}(\mathbf{z}_{i},\widehat{u}_{i,t})\|+L_{\sigma_{g,i}}\big\|[\mathbf{z}_{i,t};\widehat{u}_{i,t}]-[\tilde{\mathbf{z}}_{i,t};\tilde{\widehat{u}}_{i,t}]\big\|\bigg)+L_{g_{i}}\big\|[\mathbf{z}_{i,t};\widehat{u}_{i,t}]-[\tilde{\mathbf{z}}_{i,t};\tilde{\widehat{u}}_{i,t}]\big\|$
$=2\beta_{t}\sigma_{g,i,t-1}(\mathbf{z}_{i},\widehat{u}_{i,t})+(2\beta_{t}L_{\sigma_{g,i}}+L_{g_{i}})\big\|[\mathbf{z}_{i,t};\widehat{u}_{i,t}]-[\tilde{\mathbf{z}}_{i,t};\tilde{\widehat{u}}_{i,t}]\big\|$
$\leq 2\beta_{t}\sigma_{g,i,t-1}(\mathbf{z}_{i},\widehat{u}_{i,t})+(2\beta_{t}L_{\sigma_{g,i}}+L_{g_{i}})\|\mathbf{z}_{i,t}-\tilde{\mathbf{z}}_{i,t}\|+(2\beta_{t}L_{\sigma_{g,i}}+L_{g_{i}})\|\widehat{u}_{i,t}-\tilde{\widehat{u}}_{i,t}\|$
$\overset{\zeta_{3}}{\leq}2\beta_{t}\sigma_{g,i,t-1}(\mathbf{z}_{i},\widehat{u}_{i,t})+(2\beta_{t}L_{\sigma_{g,i}}+L_{g_{i}})\|\mathbf{z}_{i,t}-\tilde{\mathbf{z}}_{i,t}\|+2\beta_{t}(2\beta_{t}L_{\sigma_{g,i}}+L_{g_{i}})\sigma_{\widehat{u}_{i,t-1}}$
$=2\beta_{t}\sigma_{g,i,t-1}(\mathbf{z}_{i},\widehat{u}_{i,t})+2\beta_{t}(2\beta_{t}L_{\sigma_{g,i}}+L_{g_{i}})\sigma_{\widehat{u}_{i,t-1}}+(2\beta_{t}L_{\sigma_{g,i}}+L_{g_{i}})\sum_{j\in\mathbf{pa}(i)}\|\mathbf{z}_{j,t}-\tilde{\mathbf{z}}_{j,t}\|$
$\leq 2\beta_{t}\sigma_{g,i,t-1}(\mathbf{z}_{i},\widehat{u}_{i,t})+2\beta_{t}(2\beta_{t}L_{\sigma_{g}}+L_{g})\sigma_{\widehat{u}_{i,t-1}}+(2\beta_{t}L_{\sigma_{g}}+L_{g})\sum_{j\in\mathbf{pa}(i)}\|x_{j,t}-\tilde{x}_{j,t}\|$
$\overset{\zeta_{4}}{\leq}2\beta_{t}\sigma_{g,i,t-1}(\mathbf{z}_{i},\widehat{u}_{i,t})+2\beta_{t}(2\beta_{t}L_{\sigma_{g}}+L_{g})\sigma_{\widehat{u}_{i,t-1}}$
$~~~~+(2\beta_{t}L_{\sigma_{g}}+L_{g})\sum_{j\in\mathbf{pa}(i)}2\beta_{t}M^{N_{j}}(2\beta_{t}L_{\sigma_{g}}+L_{g})^{N_{j}}\sum_{h=0}^{j}\big(\sigma_{g,h,t-1}(\mathbf{z}_{h,t})+\sigma_{\widehat{u}_{h,t-1}}\big)$
$\leq 2\beta_{t}M^{N_{i}}(2\beta_{t}L_{\sigma_{g}}+L_{g})^{N_{i}}\sum_{j=0}^{i}\big(\sigma_{g,j,t-1}(\mathbf{z}_{j,t})+\sigma_{\widehat{u}_{j,t-1}}\big)$
In steps $\zeta_{1}$ and $\zeta_{2}$, we rely on the calibrated uncertainty and Lipschitz dynamics; in step $\zeta_{2}$, we also apply the triangle inequality; step $\zeta_{3}$ is by Lemma B.1; $\zeta_{4}$ applies the inductive hypothesis. ∎
Theorem 6.1 Consider the optimization problem in (4), with the SCM satisfying Assumptions B.1B.3, where $\mathcal{G}$ is known but $\mathbf{F}$ is unknown. Then with probability at least $1-\alpha$, the cumulative regret of Algorithm 1 is bounded by
$R_{T}$ $\leq\mathcal{O}(L_{T}M^{N}d\sqrt{T\gamma_{T}}).$
Proof.
The cumulative regret is
$R_{T}=\sum_{t=1}^{T}\bigg[\mathbb{E}[y|\mathbf{a}^{*}]-\mathbb{E}[y|a_{:,t}]\bigg].$
At step $t$, the instantaneous regret is $r_{t}$. By applying Lemma B.2, $r_{t}$ is bounded by
$r_{t}$ $=\mathbb{E}[y|\mathbf{F},\mathbf{a}^{*}]-\mathbb{E}[y|\mathbf{F},a_{:,t}]$
$\leq\mathbb{E}[y_{t}|\tilde{\mathbf{F}},a_{:,t}]-\mathbb{E}[y_{t}|\mathbf{F},a_{:,t}]$
$=\mathbb{E}[\|x_{i,t}-\tilde{x}_{i,t}\||a_{:,t}]$
$\leq 2\beta_{t}M^{N}(2\beta_{t}L_{\sigma_{g}}+L_{g})^{N}\mathbb{E}\bigg[\sum_{i=0}^{d}\|\sigma_{g,i,t-1}(\mathbf{z}_{i,t})\|+\|\sigma_{\widehat{u}_{i,t-1}}\|\bigg]$
Here $L_{t}=2\beta_{t}(2\beta_{t}L_{\sigma_{g}}+L_{g})^{N}$. Thus,
$r^{2}_{t}$ $\leq L_{t}^{2}M^{2N}\bigg(\mathbb{E}\bigg[\sum_{i=0}^{d}\|\sigma_{g,i,t-1}(\mathbf{z}_{i,t})\|+\|\sigma_{\widehat{u}_{i,t-1}}\|\bigg]\bigg)^{2}$
$\leq 2dL_{t}^{2}M^{2N}\mathbb{E}\bigg[\sum_{i=0}^{d}\|\sigma_{g,i,t-1}(\mathbf{z}_{i,t})\|_{2}^{2}+\|\sigma_{\widehat{u}_{i,t-1}}\|_{2}^{2}\bigg]$
We define $R^{2}_{T}$ as
$R^{2}_{T}$ $=(\sum_{t=1}^{T}r_{t})^{2}\leq T\sum_{t=1}^{T}r_{t}^{2}$
$\leq 2dTL_{T}^{2}M^{2N}\sum_{t=1}^{T}\mathbb{E}\bigg[\sum_{i=0}^{d}\|\sigma_{g,i,t-1}(\mathbf{z}_{i,t})\|_{2}^{2}+\|\sigma_{\widehat{u}_{i,t-1}}\|_{2}^{2}\bigg]$
$=2dTL_{T}^{2}M^{2N}\Gamma_{T}.$
Here,
$\Gamma_{T}=$ $\max_{(\mathbf{z},a,\widehat{u})\in\mathcal{Z}\times\mathcal{A}\times\widehat{\mathcal{U}}}\sum_{t=1}^{T}\sum_{i=0}^{d}\bigg[\|\sigma_{i,t-1}(\mathbf{z}_{i,t},a_{i,t})\|_{2}^{2}+\|\sigma_{\widehat{u}_{i,t-1}}\|_{2}^{2}\bigg]$
$\leq$ $\max_{A,\widehat{U}}\sum_{t=1}^{T}\sum_{i=0}^{d}\bigg[\|\sigma_{i,t-1}(\mathbf{z}_{i,t},a_{i,t})\|_{2}^{2}+\|\sigma_{\widehat{u}_{i,t-1}}\|_{2}^{2}\bigg]$
$\leq$ $\sum_{i=0}^{d}\max_{A_{i},\widehat{U}_{i}}\sum_{t=1}^{T}\bigg[\|\sigma_{i,t-1}(\mathbf{z}_{i,t},a_{i,t})\|_{2}^{2}+\|\sigma_{\widehat{u}_{i,t-1}}\|_{2}^{2}\bigg]$
$\leq$ $\sum_{i=0}^{d}\max_{A_{i},\widehat{U}_{i}}\sum_{t=1}^{T}\bigg[\sum_{l=1}^{d_{i}}\|\sigma_{i,t-1}(\mathbf{z}_{i,t},a_{i,t},l)\|_{2}^{2}+\|\sigma_{\widehat{u}_{i,t-1}}\|_{2}^{2}\bigg]$
$\overset{\zeta_{1}}{\leq}$ $\sum_{i=0}^{d}\frac{2}{\ln(1+\rho_{i}^{-2})}\gamma_{i,T}$
$=$ $\mathcal{O}(d\gamma_{T}).$
Here $\zeta_{1}$ is due to the upper bound of the information gain [20], and $\gamma_{T}$ will often scale sublinearly in $T$ [22]. Therefore,
$R^{2}_{T}$ $\leq 2TL_{T}^{2}M^{2N}d\mathcal{O}(d\gamma_{T}).$
And,
$R_{T}$ $\leq\mathcal{O}(L_{T}M^{N}d\sqrt{T\gamma_{T}}).$
This completes the proof of the theorem.
Appendix C Additional Experimental Analysis on EXCBO
This section presents an additional empirical analysis of the proposed EXCBO algorithm.
(a) $t=0,|\mathcal{D}_{0}|=30$                                                      (b) $t=30,|\mathcal{D}_{30}|=60$
Refer to caption    Refer to caption
(c) $Y=f(X)+\epsilon$
Figure 12: Surrogate model of EXCBO on Dropwave. The blue curve is the surrogate mean, and the blue shadow stands for the scaled standard derivation of the surrogate model; red points are samples from the ground truth model. (a) At step $t=0$, the model is initialized with 30 samples; (b) At step $t=30$, more observations are around the optimal location; (c) Samples from the ground truth function of $Y$.
Dropwave Dataset
The graph structure of the Dropwave Dataset is given by Figure 6. There are two endogenous nodes in Dropwave, i.e., $X$ and the target node $Y$. There are two action nodes associated with $X$, $a_{0}$ and $a_{1}$. Here $a_{0},a_{1}\in[0,1]$, $X=\sqrt{(c_{1}\cdot a_{0}-c_{2})^{2}+(c_{1}\cdot a_{1}-c_{2})^{2}}+\epsilon_{X}$, $c_{1}=10.24,c_{2}=5.12$, and $Y=(1.0+\cos(12.0X))/(2.0+0.5X^{2})+\epsilon_{Y}$, $\epsilon_{X},\epsilon_{Y}\sim\mathcal{N}(0,\sigma)$. We change the value of $\sigma$ to generate data samples with different noise levels.
We performed a further investigation on the surrogate model learned by the EXCBO algorithm using the Dropwave dataset at $\sigma=0.1$. The algorithm is initialized with 30 data samples, and then we plot the learned surrogate function from $X$ and $Y$ at step $t=30$. Figure 12 shows the surrogate model mean and scaled standard derivation at $t=0$ and $t=30$. Plot (c) in Figure 12 shows the data points sampled from the ground truth model, and it indicates that the dataset is very noisy. As shown in the plots, at $t=0$ data points are evenly sampled in the value span of $X$. At $t=30$, we can see that the probabilistic surrogate model drives the EXCBO algorithm to obtain more samples around the optimal location. We further notice that the variance of the surrogate decreases at the locations with dense data samples. Plots (a) and (b) show that the surrogate model employed by EXCBO improves the learning of the unknown function as well as the estimation of uncertainty along with more observations.
Rosenbrock Dataset
Figure 13 shows the graph structure of the Rosenbrock dataset. Figure 14 shows the average reward values attained by the proposed EXCBO algorithm with nine different random seeds. The surrogate model is initialized with 20 samples. As shown in the figure, similar to Dropwave the reward values increase with more observational data points in the training set. However, different from the result of Dropwave, the variance of the reward value decreases significantly with more training data samples. Another difference from Dropwave is that the performance of the EXCBO algorithm is barely affected by the noise level in the dataset.
Figure 13: Graph structure of Rosenbrock dataset.
Figure 14: Reward values achieved by EXCBO in different rounds on Rosenbrock. The results are on three different Gaussian noise levels with $\sigma=\{0.1,0.2,0.3\}$.